OpenStudy (anonymous):

A race car is moving with a velocity of 144 kilometers/hour. The driver applies the brakes, and the car comes to a halt in 12.0 seconds. What is the acceleration of the car after 12.0 seconds? a) -12.0 meters/second^2 b) -3.3 meters/second^2 c) 3.3 meters/second^2 d) 10.3 meters/second^2

OpenStudy (theeric):

This question is worded badly.. But, it wants the average acceleration while breaking. Do you have any ideas about what you'll need to be doing in this problem?

OpenStudy (anonymous):

is it distance subtracted by time, then divided by the time?

OpenStudy (theeric):

Well, you don't have the distance there... You don't need it anyway! But I'm glad you're thinking! You need to know what acceleration is. Would you like me to tell you, or do you want to tell me what it is?

OpenStudy (anonymous):

i divided 144 by the 12.0 and got 12 as my answer, is that correct?

OpenStudy (theeric):

Nope, and I'll tell you why! First of all, you have the right idea. Acceleration is the change in velocity per the change in time. But, \(\dfrac{144\ [km\ /\ h]}{12\ [s]}=12\ [km\ s\ /\ h]\). Units is the problem, there. There is also the issue of direction! The change in velocity is the final amount minus the initial amount. That is \(0-144=-144\). Think, that really is the change. See, you're at 144km/h. How do you get to 0? You change -144 to get to 0.

OpenStudy (theeric):

So, let's fix that first problem. Change the units! Do you have experience with converting things?

OpenStudy (anonymous):

not really this isnt my homework

OpenStudy (theeric):

You're asking this for someone else, you mean?

OpenStudy (anonymous):

yes

OpenStudy (theeric):

If so, here's what you can show the person who needs to do it. \(\bar a=\dfrac{v_f-v_i}{\Delta t}\) \(1\ [km]=1,000\ [m]\implies 1=\dfrac{1,000\ [m]}{1\ [km]}\) \(1\ [hr]=3,600\ [s]\implies 1= \dfrac{1\ [hr]}{3,600\ [s]}\) \(144\ [km/hr]\times1\times1\\=144\ [km/hr]\times\dfrac{1,000\ [m]}{1\ [km]}\times \dfrac{1\ [hr]}{3,600\ [s]}=40\ [m/s]\) So, your change is \(-40\ [m/s]\), and the time interval is \(12s\). Now you can use division like you tried before. That is because \(\bar a=\dfrac{v_f-v_i}{\Delta t}\).